OpenStudy (anonymous):

Hey guys I need help. Please help 4+6 (x-1)=6- 2x Can you help please?

OpenStudy (saifoo.khan):

4 + 6x - 6 = 6 -2x

OpenStudy (anonymous):

Idk

OpenStudy (saifoo.khan):

Adding like terms: 6x - 2 = 6 -2x 6x + 2x = 6 + 2 8x = 8 x = 8/8 x = 1

OpenStudy (anonymous):

agree ^^^^

OpenStudy (anonymous):

I was just guessing it was one or two.

OpenStudy (anonymous):

you have good intuition sunnyd48 !

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Thank you @ Saifoo khan :D

OpenStudy (anonymous):

Ur welcome

OpenStudy (anonymous):

and you guys too

OpenStudy (saifoo.khan):

you erlvomr.

OpenStudy (saifoo.khan):

welcome*